Express the equation in logarithmic form. (a) ex + 2 = 0.9 (b) €0.3x =
Answer #1

D 2+2 e = 0.9 Apply log on both sides log (ex+2) = log (0.9) If loga (xb) = b. loga (x) & log (ex+2) = (x+2) en (e) = (x+2) eApply log on both sides dog (60.3x) = log (t). If losa (xh) = b. loga(x). log (€0.3) = 0.34 logle) = 0.34 log (e) = log (+).